OpenStudy (anonymous):

Resolution of 300g at 60 degrees Fx'= Fy'= Help!!

OpenStudy (anonymous):

Unclear question.

OpenStudy (anonymous):

Its apart of my lab and I have no idea what to do

OpenStudy (anonymous):

your question is not clear but I try to give idea. is there some acc given? if given then Fx=ma m=0.3kg if velocity all ready given in x direction then Fx=v and Fy=Vi+at a=F/m t=d/v if y component is initially zero than Vi=0 Fy=(F/m)(d/v) tan theta = Fy/Fx